OpenStudy (anonymous):

The equation of line CD is y = -2x - 2. Write an equation of a line parallel to line CD in slope-intercept form that contains point (4, 5).

OpenStudy (anonymous):

Select one: a. y = -2x + 13 b.y=-1/2x +7 c.y=-1/2x+3 d.y=2x-3

OpenStudy (anonymous):

@Nurali @ganeshie8

OpenStudy (nurali):

you are given the equation of the line: y = -2x - 2 to get a line that is parallel to it, we take the slope of the CD line, and put it into the equation we are going to make. The slope of the CD line is the coefficient infront of the x term, so slope, m, =-2 using the slope intercept format, we can plug in what we know for our new line. y =mx+b 5 = -2(4) + b, b=13 so equation is y = -2x + 13

OpenStudy (anonymous):

thhnx